OpenStudy (anonymous):

40 is what percent of 50?

OpenStudy (abstone23):

80%

OpenStudy (gottennis121):

No direct answers please. :)

OpenStudy (anonymous):

You need to ask ypurself 50 times what is 40? 50 times 0.80 equals 40 ok

OpenStudy (anonymous):

(X percentage) * 40 = 50. i.e X percentage = 50/40 X percentage = 1.25 i.e X = 1.25 * 100 X = 125 % So 125 % of 40 is 50

OpenStudy (freckles):

40 is what percent of 50 is replace with equals what percent replace with x x is what we are trying to find 40=x50 50x=40 Divide both sides by 50. Then write x as a percent.